# Transition to inverse cascade in turbulent rotating convection in absence of the large-scale vortex

### v1

@@ 4 Transition towards inverse cascade
These results are depicted in figure 5. As first observed by Guervilly et al. (2014), when the LSV emerges, the heat transfer slightly decreases. Without hypoviscosity, however, the Nu for the case without LSV can only be measured in the transient phase before the LSV emerges, at contrast with the statistically steady state obtained here. Comparing the runs with and without hypoviscosity, we find that the suppression of Nu by the LSV remains in the order of 3-8% throughout the full range of Ra in which the LSV is active. In the cases where there is no LSV, there is a good agreement between the runs with and without hypoviscosity.
Figure 5: (a) The average Nusselt number Nu as a function of Ra for $\textrm{Ek}=10^{-4}$ and $\textrm{Pr}=1$ for the series of runs with and without hypoviscosity. (b) The same data compensated by the Nusselt number of the runs with hypoviscosity. This shows that the LSV that is present in the runs without hypoviscosity lowers the Nu by around 3-8%.
@@ 5 Conclusions and outlook
The continuous transition towards the inverse cascade that is found in this work is in agreement with what is found in other quasi-2D systems, such as magneto-hydrodynamics (Seshasayanan et al., 2014), thin layer turbulence (Benavides & Alexakis, 2017) and forced rotating turbulence (Pestana & Hickel, 2019; van Kan & Alexakis, 2020). There, indications for criticality of this transition were found in the limit of increasingly large domain size and high Reynolds number. For rotating convection, this will be hard to prove numerically, given the imperfect scale separation that is inherent to the natural convective forcing. Nonetheless, one may speculate that criticality could manifest in a similar fashion in rotating convection, considering the found morphological and theoretical similarities between rotating convection and the other quasi-2D systems.
@@
[…]
[Acknowledgments]The author thanks R.P.J. Kunnen and H.J.H. Clercx for useful discussions.
[…]
[Funding]This work is supported by the Netherlands Organization for Scientific Research (NWO) through the use of supercomputer facilities (Snellius) under Grant No. 2023.026, EINF-11692, EINF-12328. This publication is part of the project “Shaping turbulence with smart particles” with Project No. OCENW.GROOT.2019.031 of the research program Open Competitie ENW XL which is (partly) financed by the Dutch Research Council (NWO).
[…]
[Declaration of Interests]The author reports no conflict of interest.
Appendix A Validation of varying hypoviscosity
When the large scales at which the hypoviscosity is active are sufficiently well separated from the intermediate scales that dominate the energy injection, the rate at which the hypoviscosity dissipates kinetic energy becomes precisely equal to the inverse flux of kinetic energy, independent of the exact parameters of the hypodissipation. This is analogous to the relation between energy injection rate and viscous dissipation rate in normal 3D turbulence.
@@ Appendix A Validation of varying hypoviscosity
We showcase this for an exemplary run in figure 6. Indeed, for the run with higher hypoviscosity coefficient $\nu_{\alpha}=0.076$, the spectral energy at the large scales naturally becomes smaller, such that the total hypodissipation becomes equal to the run with smaller hypoviscosity coefficient $\nu_{\alpha}=0.038$, with both agreeing within error bars. This confirms that here, the hypodissipation is indeed well tuned and it is a good proxy for the total inverse flux.

### v2

@@ 4 Transition towards inverse cascade
These results are depicted in figure 5. As first observed by Guervilly et al. (2014), when the LSV emerges, the heat transfer slightly decreases. Without hypoviscosity, however, the Nu for the case without LSV can only be measured in the transient phase before the LSV emerges, at contrast with the statistically steady state obtained here. Comparing the runs with and without hypoviscosity, we find that the suppression of Nu by the LSV remains in the order of 3-8% throughout the full range of Ra in which the LSV is active for the $\textrm{Ek}=10^{-4}$ considered here. In the cases where there is no LSV, there is a good agreement between the runs with and without hypoviscosity.
It should be noted that the current simulations are conducted with stress-free boundary conditions. When considering no-slip boundaries, the effect of Ekman pumping is known to increase Nu (Stellmach et al., 2014), which could counteract the suppressive effect of the LSV.
Figure 5: (a) The average Nusselt number Nu as a function of Ra for $\textrm{Ek}=10^{-4}$ and $\textrm{Pr}=1$ for the series of runs with and without hypoviscosity and $L/H=2.235$. (b) The same data compensated by the Nu of the runs with hypoviscosity. For the runs without hypoviscosity, black diamonds denote cases that show LSV formation. This indicates that the LSV lowers the Nu by around 3-8%.
Oliver et al. (2023) have studied the scaling of Nu in absence of the LSV in the limit of $\textrm{Ek}\to 0$ using asymptotically reduced equations. Remarkably, they find that in that case, the Nu approaches the diffusivity-free scaling $\textrm{Nu}\sim\textrm{Pr}^{-1/2}\textrm{Ra}^{3/2}\textrm{Ek}^{2}$. In the current work, the Ek is not sufficiently low to retrieve such scaling.
@@ 5 Conclusions and outlook
The continuous transition towards the inverse cascade that is found in this work is in agreement with what is found in other quasi-2D systems, such as magneto-hydrodynamics (Seshasayanan et al., 2014), thin layer turbulence (Benavides & Alexakis, 2017) and forced rotating turbulence (Pestana & Hickel, 2019; van Kan & Alexakis, 2020). There, indications for criticality of this transition were found in the limit of increasingly large domain size and high Reynolds number. For rotating convection, this will be hard to prove numerically, given the imperfect scale separation that is inherent to the natural convective forcing. Nonetheless, one may speculate that criticality could manifest in a similar fashion in rotating convection, considering the found phenomenological and theoretical similarities between rotating convection and the other quasi-2D systems.
@@
{bmhead}
[Acknowledgments.] The author thanks R.P.J. Kunnen, H.J.H. Clercx and A. Alexakis for useful discussions.
{bmhead}
[Funding.] This work is supported by the Netherlands Organization for Scientific Research (NWO) through the use of supercomputer facilities (Snellius) under Grant No. 2023.026, EINF-11692, EINF-12328. This publication is part of the project “Shaping turbulence with smart particles” with Project No. OCENW.GROOT.2019.031 of the research program Open Competitie ENW XL which is (partly) financed by the Dutch Research Council (NWO).
{bmhead}
[Declaration of Interests.] The author reports no conflict of interest.
{appen}
6 Other methods of LSV suppression
In this work, an adapted form of hypoviscosity is used to suppress the formation of the LSV. In practice, however, different mechanisms could be responsible for the suppression of the LSV. A very natural mechanism could be the friction caused by no-slip boundaries, which give rise to Ekman boundary layers that can partially or fully suppress the formation of the LSV (Kunnen et al., 2016; Julien et al., 2016; Plumley et al., 2016; Aguirre Guzmán et al., 2020). Another approach for suppression of the LSV is put forward by Oliver et al. (2023), where the non-linear term of the Navier-Stokes equations is modified to remove the non-linear interactions in the 2D vertically averaged manifold, modifying equation (2a) as
$\frac{\partial\bm{u}}{\partial t}+\left(\bm{u}\cdot\bm{\nabla}\right)\bm{u}-\mkern 1.5mu\overline{\mkern-1.5mu\left(\bm{u}\cdot\bm{\nabla}\right)\bm{u}\mkern-1.5mu}\mkern 1.5mu+\frac{\left(\textrm{Pr}/\textrm{Ra}\right)^{1/2}}{\textrm{Ek}}\bm{e}_{z}\times\bm{u}=-\bm{\nabla}p+T\bm{e}_{z}+\left(\frac{\textrm{Pr}}{\textrm{Ra}}\right)^{1/2}\Delta\bm{u}.$ (10)
In figure 6, we compare the different methods of LSV suppression for an exemplary case. It shows that for the case considered here, all mechanisms yield a complete suppression of the LSV formation at large scales, while other scales remain mostly unaffected. The simulation with hypoviscosity yields the least aggressive suppression of the large scale energy. We finally point out that in our work, the hypoviscosity not only serves to suppress the formation of the LSV, but also serves as a tool to measure the strength of the inverse cascade. How to harness other mechanisms in order to measure the strength of the inverse cascade may be less straightforward.
Figure 6: Comparison of different methods of LSV suppression for an exemplary case with $\textrm{Ra}=8\times 10^{6}$ and $L/H=2.235$, depicted through instantaneous kinetic energy spectra (a) and snapshots of kinetic energy at the mid-plane (b-e). We consider the case with LSV, without suppression (b), the case with hypoviscosity as treated in the main text (c), the case with no-slip top and bottom wall boundary conditions (d) and the case with suppression of 2D non-linear interactions (e). For comparison, we also show the case with LSV where the $k=2\pi/L$ modes have been fully filtered out a posteriori (f).
7 Validation with varying hypoviscosity
When the large scales at which the hypoviscosity is active are sufficiently well separated from the intermediate scales that dominate the energy injection, the rate at which the hypoviscosity dissipates kinetic energy becomes precisely equal to the inverse flux of kinetic energy, independent of the exact parameters of the hypodissipation. This is analogous to the relation between energy injection rate and viscous dissipation rate in regular 3D turbulence.
We show this for an exemplary case in figure 7. Indeed, for the run with higher hypoviscosity coefficient $\nu_{\alpha}=0.076$, the spectral energy at the large scales naturally becomes smaller, such that the total hypodissipation becomes equal to the run with smaller hypoviscosity coefficient $\nu_{\alpha}=0.038$, with both agreeing within error bars. This confirms that here, the hypodissipation is indeed well tuned and it is a good proxy for the total inverse flux.
@@ 7 Validation with varying hypoviscosity
8 Validation of absence of hysteresis
In earlier works, it has been observed that the transition into the LSV state at high Ra is discontinuous and hysteretic (Favier et al., 2019; de Wit et al., 2022a). In the current work, we show that when the LSV is suppressed using a friction mechanism at large scales, the transition towards the inverse cascade regime becomes continuous. As a result, we expect that also the hysteresis in the transition should vanish. In figure 8, we test this explicitly for an exemplary case with $\textrm{Ra}=2.5\times 10^{7}$, which is in the range where hysteresis was observed for the case with LSV (de Wit et al., 2022a). Indeed, comparing the run that is freshly initialized as described in the main text, to the runs that are restarted from a lower $\textrm{Ra}=1.0\times 10^{7}$ and a higher $\textrm{Ra}=4.0\times 10^{7}$, we confirm that there is no statistical dependence on the initial conditions.
Figure 8: Timeseries of hypodissipation for the case with $\textrm{Ra}=1.5\times 10^{7}$ and $L/H=4.470$ for a run that is freshly initialized (green line), a run that is restarted from a run with $\textrm{Ra}=1.0\times 10^{7}$ (purple line) and a run that is restarted from a run with $\textrm{Ra}=4.0\times 10^{7}$ (light blue line). The gray shaded area indicates the time interval in which the statistically steady state is reached that is used for averaging, while the horizontal dashed lines indicate the averages themselves, with statistical errorbars as coloured shaded areas.